OpenStudy (anonymous):

Solve 2x + 4(x – 6) = 3x + 9

OpenStudy (anonymous):

2x+4x-24=3x+9 6x-24=3x+9 6x-3x=9+24 3x=33 x=33/3 x=11

OpenStudy (anonymous):

Thank you